# FedAPTA: Federated Multi-task Learning for Heterogeneous Devices with Adaptive Layer-wise Pruning and Task-aware Aggregation

### v2

@@ 2.1 Overview
In this study, we propose FedAPTA, a FL framework specifically designed to efficiently train task-specific models across heterogeneous devices, where different devices have different computing resources and may undertake different tasks. Unlike traditional FL frameworks, which focus on training and maintaining a shared model structure for all devices, FedAPTA allows each device to possess a local model customized for its computational capability. The overall operational workflow of FedAPTA is illustrated in Fig. 1, which includes two fundamental steps: the training of local models on devices and the aggregation of models on the central server. Particularly, the central server uses a pre-trained model as the initialized global model $w$ and distributes it to all devices. $w$ can come from pre-trained models trained on openly accessible datasets. To facilitate efficient training in scenarios with heterogeneous client computational capabilities, we introduce an adaptive layer-wise model pruning method, which will be described in detail in Section $2.$ In addition, regarding to the model aggregation process, there are two central challenges that must be addressed: on one hand, to aggregate the pruned heterogeneous models, we utilize information from the latest global model to restore the complete structure of the local models, facilitating the aggregation of the pruned models (as described in Section 2.3); on the other hand, to aggregate different task-specific models, we cluster all devices by their local models to distinguish their tasks (as elaborated in Section 2.4). The complete implementation process of the proposed method is presented in Algorithm 1, and the key symbols used in this paper are listed in Table 1.
@@ 2.2 Adaptive Layer-wise Model Pruning
To begin with, for the model deployed on each device $i$, FedAPTA evaluates the importance score of each candidate layer $k$ within the model $w_{i}$ by employing the L1 norm [liu2017learning_pr1], obtaining the importance score vector $I=[I_{1},\cdots,I_{k},\cdots,I_{n}]$. Subsequently, while ensuring that the overall pruning ratio of the model remains consistent with the predefined whole pruning ratio target $\rho_{i}$, the pruning ratio $\rho_{k}$ for each candidate layer $k$ is assigned based on its parameter quantity and importance score. Specifically, let the parameter count of each candidate layer $k$ be represented by the vector $N=[N_{1},\cdots,N_{k},\cdots,N_{n}]$; this vector is then reordered according to the descending order of II. The objective is to assign a pruning ratio $\rho_{k}$ to each candidate layer $k$, where $\rho_{k}$ is between 0 and 1, such that the aggregate pruning across all layers adheres to the specified overall pruning ratio $\rho_{i}$, while the sequence of assigned pruning ratios exhibits a non-decreasing structure, it means that the layers that are deemed less important are assigned a pruning ratio that is no less than the ratios of more important layers. Let $P_{i}=[\rho_{1},\cdots,\rho_{k},\cdots,\rho_{n}]$ represent the optimization variable, then the formulation of the optimization problem is given as follows:
@@ 2.3 Heterogeneous Model Recovery
where $\hat{w_{i}}$ is the model to be recovered and $w_{t}$ is the latest global model of task $t$, which is executed by device ii. The term $(1-M_{i})$ effectively identifies the pruned parameters of local models, then the central server replaces these pruned values with their corresponding entries from the latest global model.

### v3

@@ 2.1 Overview
In this study, we propose FedAPTA, a FL framework specifically designed to efficiently train task-specific models across heterogeneous devices, where different devices have different computing resources and may undertake different tasks. Unlike traditional FL frameworks, which focus on training and maintaining a shared model structure for all devices, FedAPTA allows each device to possess a local model customized for its computational capability. The overall operational workflow of FedAPTA is illustrated in Fig. 1, which includes two fundamental steps: the training of local models on devices and the aggregation of models on the central server. Particularly, the central server uses a pre-trained model as the initialized global model $w$ and distributes it to all devices. $w$ can come from pre-trained models trained on openly accessible datasets. To facilitate efficient training in scenarios with heterogeneous client computational capabilities, we introduce an adaptive layer-wise model pruning method, which will be described in detail in Section 2.2. In addition, regarding to the model aggregation process, there are two central challenges that must be addressed: on one hand, to aggregate the pruned heterogeneous models, we utilize information from the latest global model to restore the complete structure of the local models, facilitating the aggregation of the pruned models (as described in Section 2.3); on the other hand, to aggregate different task-specific models, we cluster all devices by their local models to distinguish their tasks (as elaborated in Section 2.4). The complete implementation process of the proposed method is presented in Algorithm 1, and the key symbols used in this paper are listed in Table 1.
@@ 2.2 Adaptive Layer-wise Model Pruning
To begin with, for the model deployed on each device $i$, FedAPTA evaluates the importance score of each candidate layer $k$ within the model $w_{i}$ by employing the L1 norm [liu2017learning_pr1], obtaining the importance score vector $I=[I_{1},\cdots,I_{k},\cdots,I_{n}]$. Subsequently, while ensuring that the overall pruning ratio of the model remains consistent with the predefined whole pruning ratio target $\rho_{i}$, the pruning ratio $\rho_{k}$ for each candidate layer $k$ is assigned based on its parameter quantity and importance score. Specifically, let the parameter count of each candidate layer $k$ be represented by the vector $N=[N_{1},\cdots,N_{k},\cdots,N_{n}]$; this vector is then reordered according to the descending order of $I$. The objective is to assign a pruning ratio $\rho_{k}$ to each candidate layer $k$, where $\rho_{k}$ is between 0 and 1, such that the aggregate pruning across all layers adheres to the specified overall pruning ratio $\rho_{i}$, while the sequence of assigned pruning ratios exhibits a non-decreasing structure, it means that the layers that are deemed less important are assigned a pruning ratio that is no less than the ratios of more important layers. Let $P_{i}=[\rho_{1},\cdots,\rho_{k},\cdots,\rho_{n}]$ represent the optimization variable, then the formulation of the optimization problem is given as follows:
@@ 2.3 Heterogeneous Model Recovery
where $\hat{w_{i}}$ is the model to be recovered and $w_{t}$ is the latest global model of task $t$, which is executed by device $i$. The term $(1-M_{i})$ effectively identifies the pruned parameters of local models, then the central server replaces these pruned values with their corresponding entries from the latest global model.